Please help me. I can't answer these problems
1. (-5x³y⁴)(-5x³yz)
2. (x^9 - y^5)(x^9 + y^5)
3. (a - b)(a² - ab + b²)
4. (4x - 3y)(16x² - 12xy + 9y²)
5. (2a - 7y)²

Answer by Boreal(15235)

1. (-5x³y⁴)(-5x³yz)=25x^6y^5z Multiply the constants first, then the x terms, adding the exponents, then the y terms, and then the z terms.
2. (x^9 - y^5)(x^9 + y^5) FOIL x^18 (first) +x^9y^5 (outer)-x^9y^5 (inner)-y^10 (last)= x^18-y^10, since the second and third terms cancel each other.
3. (a - b)(a² - ab + b²) first and second terms by each of the 3 terms; there are 6 altogether:
a^3-a^2b+ab^2-a^2b+ab^2-b^3=a^3-2a^2b+2ab^2-b^3. Look for like terms, and there are two terms, a^2b and ab^2, that have two.
4. (4x - 3y)(16x² - 12xy + 9y²)=20x^3-48x^2y+36xy^2-48x^2y+36xy^2-27y^3=
20x^3-96x^2y+72xy^2-27y^3
5. (2a - 7y)²=
(2a-7y)(2a-7y) and FOIL
OR
square the first, twice the internal product, square the second
4a^2-2(14ay)+49y^2=
4a^2-28ay+49y^2